a beach ball has a diameter of 15 cm. what is the surface of the beach ball if pi=3.14
Answer by stanbon(75887)

a beach ball has a diameter of 15 cm. what is the
surface of the beach ball if pi=3.14
----
Surface area = 4(pi)r^2
---
S.A. = 4(3.14)(7.5)^2 = 706.5 sq. cm.
====================================
Cheers,
Stan H.
=============